A researcher conducts a related measures T test and gets a value of MD = 4, with a sample size of n=11 and a T value of -3. If the researcher measures effect size using the percentage of variance accounted for, what value will be obtained for r2?

a. 9/20

b. 9/19

c. -9/19

d. it is not possible to have a negative value for T
Answer #1

Answer :

Given data is :

n = 11

t value = -3

degree of freedom df = n - 1 = 11 - 1 = 10

Here we need to find the :

Option b is correct